3. Reduce the quadratic form x 1 ^ 2 +5 x 2 ^ 2 + x 3 ^ 2 +2x 1 x 2 +2x 2 x 3 +6x 3 x 1 to canonical form through an orthogonal transformation and also find its rank, index, signature and nature of the quadratic form.
1
Expert's answer
2020-12-29T17:11:58-0500

The matrix of the quadratic form is


"A=\begin{bmatrix}
 1 & 1 & 3 \\
 1 & 5 & 1 \\
 3 & 1 & 1 \\
\end{bmatrix}"

Find the eigenvalues and eigenvectors of A


"A-\lambda I=\begin{bmatrix}
 1-\lambda & 1 & 3 \\
 1 & 5-\lambda & 1 \\
 3 & 1 & 1-\lambda \\
\end{bmatrix}"

"\begin{vmatrix}
 1-\lambda & 1 & 3 \\
 1 & 5-\lambda & 1 \\
 3 & 1 & 1-\lambda \\
\end{vmatrix}=(1-\lambda)\begin{vmatrix}
 5-\lambda & 1 \\
 1 & 1-\lambda
\end{vmatrix}"

"-(1)\begin{vmatrix}
 1 & 1 \\
 3 & 1-\lambda
\end{vmatrix}+(3)\begin{vmatrix}
 1 & 5-\lambda \\
 3 & 1
\end{vmatrix}"

"=(1-\lambda)^2(5-\lambda)-1+\lambda-1+\lambda+3+3-45+9\lambda"

"=-\lambda^3+7\lambda^2-36"

This is a characteristic polynomial.

Solve the equation


"-\lambda^3+7\lambda^2-36=0"

"-\lambda^2(\lambda-6)+(\lambda-6)(\lambda+6)=0"

"-(\lambda-6)(\lambda-3)(\lambda+2)=0"

The roots are: "\lambda_1=6,\lambda_2=3,\lambda_3=-2."

These are the eigenvalues.

Find the eigenvectors.

"\lambda_1=6"


"\begin{bmatrix}
 1-\lambda & 1 & 3 \\
 1 & 5-\lambda & 1 \\
 3 & 1 & 1-\lambda \\
\end{bmatrix}=\begin{bmatrix}
 -5 & 1 & 3 \\
 1 & -1 & 1 \\
 3 & 1 & -5 \\
\end{bmatrix}"

"R_2=R_2+(1\/5)R_1"


"\begin{bmatrix}
 -5 & 1 & 3 \\
 0 & -4/5 & 8/5 \\
 3 & 1 & -5 \\
\end{bmatrix}"

"R_3=R_3+(3\/5)R_1"


"\begin{bmatrix}
 -5 & 1 & 3 \\
 0 & -4/5 & 8/5 \\
 0 & 8/5 & -16/5 \\
\end{bmatrix}"

"R_3=R_3+(2)R_2"


"\begin{bmatrix}
 -5 & 1 & 3 \\
 0 & -4/5 & 8/5 \\
 0 & 0 & 0 \\
\end{bmatrix}"

"R_2=(-5\/4)R_2"


"\begin{bmatrix}
 -5 & 1 & 3 \\
 0 & 1 & -2 \\
 0 & 0 & 0 \\
\end{bmatrix}"

"R_1=R_1-R_2"


"\begin{bmatrix}
 -5 & 0 & 5 \\
 0 & 1 & -2 \\
 0 & 0 & 0 \\
\end{bmatrix}"

"R_1=(-1\/5)R_1"


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & -2 \\
 0 & 0 & 0 \\
\end{bmatrix}"

Solve the matrix equation


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & -2 \\
 0 & 0 & 0 \\
\end{bmatrix}\begin{bmatrix}
 v _1 \\
 v_2 \\
 v_3
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0 \\
 0
\end{bmatrix}"

If we take "v_3=t," then "v_1=t,v_2=2t,v_3=t."


"\text{v}=\begin{bmatrix}
 t \\
 2t \\
 t
\end{bmatrix}=\begin{bmatrix}
 1 \\
 2 \\
 1
\end{bmatrix}t"

"\lambda_2=3"

"\begin{bmatrix}
 1-\lambda & 1 & 3 \\
 1 & 5-\lambda & 1 \\
 3 & 1 & 1-\lambda \\
\end{bmatrix}=\begin{bmatrix}
 -2 & 1 & 3 \\
 1 & 2 & 1 \\
 3 & 1 & -2 \\
\end{bmatrix}"

"R_2=R_2+(1\/2)R_1"

"\begin{bmatrix}
 -2 & 1 & 3 \\
 0 & 5/2 & 5/2 \\
 3 & 1 & -2 \\
\end{bmatrix}"

"R_3=R_3+(3\/2)R_1"

"\begin{bmatrix}
 -2 & 1 & 3 \\
 0 & 5/2 & 5/2 \\
 0 & 5/2 & 5/2 \\
\end{bmatrix}"

"R_3=R_3-R_2"

"\begin{bmatrix}
 -2 & 1 & 3 \\
 0 & 5/2 & 5/2 \\
 0 & 0 & 0 \\
\end{bmatrix}"

"R_2=(2\/5)R_2"

"\begin{bmatrix}
 -2 & 1 & 3 \\
 0 & 1 & 1 \\
 0 & 0 & 0 \\
\end{bmatrix}"

"R_1=R_1-R_2"

"\begin{bmatrix}
 -2 & 0 & 2 \\
 0 & 1 & 1 \\
 0 & 0 & 0 \\
\end{bmatrix}"

"R_1=(-1\/2)R_1"

"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & 1 \\
 0 & 0 & 0 \\
\end{bmatrix}"

Solve the matrix equation


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & 1 \\
 0 & 0 & 0 \\
\end{bmatrix}\begin{bmatrix}
 v _1 \\
 v_2 \\
 v_3
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0 \\
 0
\end{bmatrix}"

If we take "v_3=t," then "v_1=t,v_2=-t,v_3=t."

"\text{v}=\begin{bmatrix}
 t \\
 -t \\
 t
\end{bmatrix}=\begin{bmatrix}
 1 \\
 -1 \\
 1
\end{bmatrix}t"



"\lambda_3=-2"

"\begin{bmatrix}
 1-\lambda & 1 & 3 \\
 1 & 5-\lambda & 1 \\
 3 & 1 & 1-\lambda \\
\end{bmatrix}=\begin{bmatrix}
 3 & 1 & 3 \\
 1 & 7 & 1 \\
 3 & 1 & 3 \\
\end{bmatrix}"

"R_2=R_2-(1\/3)R_1"

"\begin{bmatrix}
 3 & 1 & 3 \\
 0 & 20/3 & 0 \\
 3 & 1 & 3 \\
\end{bmatrix}"

"R_3=R_3-R_1"

"\begin{bmatrix}
 3 & 1 & 3 \\
 0 & 20/3 & 0 \\
 0 & 0 & 0 \\
\end{bmatrix}"


"R_2=(3\/20)R_2"

"\begin{bmatrix}
 3 & 1 & 3 \\
 0 & 1 & 0 \\
 0 & 0 & 0 \\
\end{bmatrix}"

"R_1=R_1-R_2"

"\begin{bmatrix}
 3 & 0 & 3 \\
 0 & 1 & 0 \\
 0 & 0 & 0 \\
\end{bmatrix}"

"R_1=(1\/3)R_1"

"\begin{bmatrix}
 1 & 0 & 1 \\
 0 & 1 & 0 \\
 0 & 0 & 0 \\
\end{bmatrix}"


Solve the matrix equation


"\begin{bmatrix}
 1 & 0 & 1 \\
 0 & 1 & 0 \\
 0 & 0 & 0 \\
\end{bmatrix}\begin{bmatrix}
 v _1 \\
 v_2 \\
 v_3
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0 \\
 0
\end{bmatrix}"

If we take "v_3=t," then "v_1=-t,v_2=0,v_3=t."

"\text{v}=\begin{bmatrix}
 -t \\
 0 \\
 t
\end{bmatrix}=\begin{bmatrix}
 -1 \\
 0 \\
 1
\end{bmatrix}t"


The Normalized matrix


"N=\begin{bmatrix}
 1/\sqrt{6} & 1/\sqrt{3} & -1/\sqrt{2} \\
 2/\sqrt{6} & -1/\sqrt{3} & 0 \\
 1/\sqrt{6} & 1/\sqrt{3} & 1/\sqrt{2} \\
\end{bmatrix}"

"N^T=\begin{bmatrix}
 1/\sqrt{6} & 2/\sqrt{6} & 1/\sqrt{6} \\
 1/\sqrt{3} & -1/\sqrt{3} & 1/\sqrt{3} \\
 -1/\sqrt{2} & 0 & 1/\sqrt{2} \\
\end{bmatrix}"

"\begin{bmatrix}
 1/\sqrt{6} & 2/\sqrt{6} & 1/\sqrt{6} \\
 1/\sqrt{3} & -1/\sqrt{3} & 1/\sqrt{3} \\
 -1/\sqrt{2} & 0 & 1/\sqrt{2} \\
\end{bmatrix}A\begin{bmatrix}
 1/\sqrt{6} & 1/\sqrt{3} & -1/\sqrt{2} \\
 2/\sqrt{6} & -1/\sqrt{3} & 0 \\
 1/\sqrt{6} & 1/\sqrt{3} & 1/\sqrt{2} \\
\end{bmatrix}="

"=\begin{bmatrix}
 6 & 0 & 0 \\
 0 & 3 & 0 \\
 0 & 0 & -2 \\
\end{bmatrix}"

"\begin{bmatrix}
 y_1 & y_2 & y_3
\end{bmatrix}\begin{bmatrix}
 6 & 0 & 0 \\
 0 & 3 & 0 \\
 0 & 0 & -2 \\
\end{bmatrix}\begin{bmatrix}
 y_1 \\
 y_2 \\
 y_3
\end{bmatrix}="

"=6y_1^2+3y_2^2-2y_3^2"

 which is the required canonical form.


"Rank=3" (the number of non-zero eigenvalues).

"Index=2" (the number of positive eigenvalues).

"Signature=2-1=1" (the difference between the number of positive eigenvalues and the number of negative eigenvalues ).

 

Nature: the eigenvalues are "6>0,3>0,-2<0." Therefore the given quadratic form is indefinite in nature.